# Integrable measure equivalence rigidity of right-angled Artin groups via quasi-isometry

### v1

@@ 4.3 Restricting couplings to stabilizers
We denote by $\kappa$ the common cardinality of all $\mathsf{G}$-stabilizers on $\mathbb{D}$. Given two elements $s,u\in\mathbb{D}$, we let $L^{0}_{s\to u}$ be the Borel subset of $L^{0}$ consisting of all elements that send $s$ to $u$, and we let $\Omega_{s\to u}=\theta^{-1}(L^{0}_{s\to u})$. Notice that $\Omega_{s\to u}$ is invariant under $\mathsf{G}_{u}$ – and on the other hand $\mu(g\Omega_{s\to u}\cap\Omega_{s\to u})=0$ for every $g\in\mathsf{G}\setminus\mathsf{G}_{u}$. In addition, the $\mathsf{G}$-translates of $\Omega_{s\to u}$ cover $\Omega$ because the $\mathsf{G}$-action on $\mathbb{D}$ is transitive. Thus, any Borel fundamental domain for the action of $\mathsf{G}_{u}$ on $\Omega_{s\to u}$ is also a fundamental domain for the action of $\mathsf{G}$ on $\Omega$. The measure of any such fundamental domain is equal to $\mu(\Omega_{s\to u})/|\mathsf{G}_{u}|=\mu(\Omega_{s\to u})/\kappa$. In particular, as $s,u$ vary in the $\mathsf{G}$-orbit of $v$, the sets $\Omega_{s\to u}$ all have the same (positive) measure, which we denote by $m$. Corollary 4.4 (applied with $L=L^{0}$ and $K=\mathbb{D}$, so that $\mathrm{Aut}(K)=\mathrm{Bij}(\mathbb{D})$) ensures that the $\mathsf{H}$-stabilizer of any element $u\in\mathbb{D}$ is measure equivalent to $\mathsf{G}_{u}$, whence finite. Let $k$ be a bound on the cardinality of a finite subgroup of $\mathsf{H}$. For every $v\in\mathsf{G}\cdot u$, the set $\Omega_{u\to v}$ is $\mathsf{H}_{u}$-invariant; we let $\Omega^{\prime}_{u\to v}\subseteq\Omega_{u\to v}$ be a fundamental domain for the action of the finite group $\mathsf{H}_{u}$. Then $\Omega^{\prime}_{u\to v}$ has measure at least $m/k$. One has $\Omega=\coprod_{u\in\mathbb{D}}\Omega_{u\to v}$. Therefore, if $T\subseteq\mathbb{D}$ is a set of representatives of the $\mathsf{H}$-orbits, then $\coprod_{u\in T}\Omega^{\prime}_{u\to v}$ is a fundamental domain for the $\mathsf{H}$-action on $\Omega$. Since $\mathsf{H}$ has a finite measure fundamental domain, and the measure of the above fundamental domain is at least equal to $m|T|/k$, it follows that $|T|<+\infty$. This concludes our proof. ∎
@@
Let $\mathsf{G}$ be a countable group, acting faithfully by automorphisms on a polyhedral complex $K$ with countably many cells, in such a way that $\mathsf{G}$ and $\mathrm{Aut}(K)$ have the same orbits of vertices. Let $\mathsf{H}$ be a countable group with bounded torsion that is measure equivalent to $\mathsf{G}$, let $(\Omega,\mu)$ be a measure equivalence coupling between $\mathsf{G}$ and $\mathsf{H}$, and assume we are given a homomorphism $\iota:\mathsf{H}\to\mathrm{Aut}(K)$ and a measurable almost $(\mathsf{G}\times\mathsf{H})$-equivariant map $\theta:\Omega\to\mathrm{Aut}(K)$.
@@
Proposition 4.6 and Corollary 4.7 fail if one does not assume that $\mathsf{H}$ has bounded torsion. Indeed, it is possible to find a countable group $\mathsf{G}$ acting properly and cocompactly on a polyhedral complex $K$, satisfying the assumptions of Corollary 4.7, and a non-uniform lattice $\mathsf{H}$ in $\mathrm{Aut}(K)$, acting with infinitely many orbits of vertices. In this case $\mathsf{G}$ and $\mathsf{H}$ are measure equivalent.
@@ 4.4 Measure equivalence couplings and index
Let $\mathsf{G}$ and $\mathsf{H}$ be two countable groups, and let $\Sigma$ be a measure equivalence coupling between $\mathsf{G}$ and $\mathsf{H}$. Let $\mathsf{H}^{\prime}\subset\mathsf{H}$ be a subgroup. Assume that there exists a positive measure $(\mathsf{G}\times\mathsf{H}^{\prime})$-invariant Borel subset $\Sigma^{\prime}\subset\Sigma$ such that for any $h\in\mathsf{H}\setminus\mathsf{H}^{\prime}$ we have $\mu(h\Sigma^{\prime}\cap\Sigma^{\prime})=0$. Then $\mathsf{H}^{\prime}$ is of finite index in $\mathsf{H}$.
@@
Let $E$ and $E^{\prime}$ be respective Borel fundamental domains for the actions of $\mathsf{G}$ and $\mathsf{H}^{\prime}$ on $\Sigma^{\prime}$, chosen such that $U=E\cap E^{\prime}$ has positive measure (these exist because $\mathsf{G}$ and $\mathsf{H}$ admit Borel fundamental domains on $\Sigma$). We claim that $E$ is contained in a Borel fundamental domain $X_{\mathsf{G}}$ for the $\mathsf{G}$-action on $\Sigma$, and $E^{\prime}$ is contained in a Borel fundamental domain $X_{\mathsf{H}}$ for the $\mathsf{H}$-action on $\Sigma$. For the first part of the claim, we take any Borel fundamental domain $X^{\prime}_{\mathsf{G}}$ for the $\mathsf{G}$-action on $\Sigma$ and take $X_{\mathsf{G}}=(X^{\prime}_{\mathsf{G}}\setminus\Sigma^{\prime})\cup E$. For the second part of the claim, it suffices to prove that for any $h\in\mathsf{H}$, the intersection $hE^{\prime}\cap E^{\prime}$ has measure zero. If $h\in\mathsf{H}^{\prime}$, then this follows from that $E^{\prime}$ is a fundamental domain for the $\mathsf{H}^{\prime}$-action on $\Sigma^{\prime}$. If $h\notin\mathsf{H}^{\prime}$, then our assumption implies that $\mu(h\Sigma^{\prime}\cap\Sigma^{\prime})=0$, and the claim follows. As recalled in Section 4.1, there is a natural measure-preserving action $\mathsf{H}\curvearrowright X_{\mathsf{G}}$, obtained from the identification $X_{\mathsf{G}}\approx\mathsf{G}\backslash\Sigma$. We let $\mathcal{G}_{1}$ be the associated measured groupoid over $X_{\mathsf{G}}$; it is equipped with a natural cocycle $\mathcal{G}_{1}\to\mathsf{H}$. Similarly, we have a natural measure-preserving action $\mathsf{G}\curvearrowright X_{\mathsf{H}}$, and we let $\mathcal{G}_{2}$ be the associated measured groupoid over $X_{\mathsf{H}}$. It is equipped with a natural cocycle $\mathcal{G}_{2}\to\mathsf{G}$. Let now $U_{1}=X_{\mathsf{G}}\cap X_{\mathsf{H}}$: this is a positive measure Borel subset (because it contains $U=E\cap E^{\prime}$). We have $(\mathcal{G}_{1})_{|U_{1}}=(\mathcal{G}_{2})_{|U_{1}}$. We denote by $\mathcal{G}$ this measured groupoid over $U_{1}$; up to restricting $U_{1}$ to a conull Borel subset, $\mathcal{G}$ is naturally equipped with two cocycles $\rho_{\mathsf{G}}:\mathcal{G}\to\mathsf{G}$ and $\rho_{\mathsf{H}}:\mathcal{G}\to\mathsf{H}$. We now claim that $\mathcal{G}_{|U}=\rho_{\mathsf{H}}^{-1}(\mathsf{H}^{\prime})_{|U}$. The lemma follows from this claim by using e.g. [HH22b, Lemma B.3]. To prove the claim, it suffices to show that $\mathcal{G}_{|U}\subseteq\rho_{\mathsf{H}}^{-1}(\mathsf{H}^{\prime})_{|U}$. Let $\gamma\in\mathcal{G}_{|U}$, let $u=s(\gamma)$ be the source of $\gamma$. As $\Sigma^{\prime}$ is $\mathsf{G}$-invariant, we have $gu\in\Sigma^{\prime}$. Since $E^{\prime}$ is a fundamental domain for the action of $\mathsf{H}^{\prime}$ on $\Sigma^{\prime}$, there exists a unique $h\in\mathsf{H}^{\prime}$ such that $hgu\in E^{\prime}$. So $h$ is also the unique element of $\mathsf{H}$ satisfying $hgu\in X_{\mathsf{H}}$. Therefore $\rho_{\mathsf{H}}(\gamma)=h$, showing that $\gamma\in\rho_{\mathsf{H}}^{-1}(\mathsf{H}^{\prime})$, as desired. ∎
@@ 5 Proximal dynamics and strong ICC property for $\mathrm{Aut}(\mathbb{B})$
The goal of this section is to prove the following proposition. We refer to the first paragraph of Section 4.2 for the definition of the strong ICC property.
@@
Let $G$ be a right-angled Artin group with trivial center, and let $\mathbb{B}$ be its right-angled building.
Then the inclusion $G\subseteq\mathrm{Aut}(\mathbb{B})$ is strongly ICC.
@@ 5.1 A strong ICC lemma, after Bader–Furman–Sauer
Recall that an action of a countable group $G$ by homeomorphisms on a compact metrizable space $K$ is strongly proximal in the sense of Furstenberg [Fur63] if the $G$-orbit of every $\nu\in\mathrm{Prob}(K)$ contains a Dirac mass in its weak-*** closure. The action is minimal if every $G$-orbit is dense in $K$. Given a compact metrizable space $K$, the group $\mathrm{Homeo}(K)$ is equipped with the topology of uniform convergence, which turns it into a Polish group. The following lemma is a small variation over an argument of Bader–Furman–Sauer [BFS13, Lemma 2.4].
@@
be the space of Borel $\mu$-stationary probability measures on $K$, which is a nonempty compact subset of $\mathrm{Prob}(K)$, equipped with the weak-*** topology. Note that $\mathrm{Prob}_{\mu}(K)$ is $G$-invariant, using that $\mu$ is conjugation-invariant.
@@
We are thus left with proving the above claim. Let $\nu\in\mathrm{Prob}_{\mu}(K)$. For every $i\in\{1,\dots,k\}$, let $\nu_{i}\in\mathrm{Prob}(K_{i})$ be the pushforward of $\nu$ under the projection map $K\to K_{i}$. Since the $G_{i}$-action on $K_{i}$ is minimal and strongly proximal, we can find a sequence $(g_{i,n})_{n\in\mathbb{N}}\in G_{i}^{\mathbb{N}}$ such that $(g_{i,n})_{\ast}\nu_{i}$ converges to $\delta_{x_{i}}$, as $n$ goes to +$\infty$+ ∞, in the weak-*** topology.
For every $n\in\mathbb{N}$, let $g_{n}=(g_{1,n},\dots,g_{k,n})$. We will now prove that the probability measures $(g_{n})_{\ast}\nu$ converge to $\delta_{x}$, as $n$ goes to +$\infty$+ ∞, in the weak-*** topology. This will conclude our proof.
@@
as $n$ goes to +$\infty$+ ∞. Intersecting these sets as $i$ varies in $\{1,\dots,k\}$, we obtain that $\nu(g_{n}^{-1}(U_{1}\times\dots\times U_{k}))\to 1$ as $n$ goes to +$\infty$+ ∞. In particular $(g_{n})_{\ast}\nu(U)\to 1$. Since $U$ was an arbitrary open set containing $x$, by the Portmanteau Theorem, this is enough to prove that $(g_{n})_{\ast}\nu$ converges to $\delta_{x}$, as desired. ∎

### v2

@@ 4.3 Restricting couplings to stabilizers
We denote by $\kappa$ the common cardinality of all $\mathsf{G}$-stabilizers on $\mathbb{D}$. Given two elements $s,u\in\mathbb{D}$, we let $L^{0}_{s\to u}$ be the Borel subset of $L^{0}$ consisting of all elements that send $s$ to $u$, and we let $\Omega_{s\to u}=\theta^{-1}(L^{0}_{s\to u})$. Notice that $\Omega_{s\to u}$ is invariant under $\mathsf{G}_{u}$ – and on the other hand $\mu(g\Omega_{s\to u}\cap\Omega_{s\to u})=0$ for every $g\in\mathsf{G}\setminus\mathsf{G}_{u}$. In addition, the $\mathsf{G}$-translates of $\Omega_{s\to u}$ cover $\Omega$ because the $\mathsf{G}$-action on $\mathbb{D}$ is transitive. Thus, any Borel fundamental domain for the action of $\mathsf{G}_{u}$ on $\Omega_{s\to u}$ is also a fundamental domain for the action of $\mathsf{G}$ on $\Omega$. The measure of any such fundamental domain is equal to $\mu(\Omega_{s\to u})/|\mathsf{G}_{u}|=\mu(\Omega_{s\to u})/\kappa$. In particular, as $s,u$ vary in the $\mathsf{G}$-orbit of $v$, the sets $\Omega_{s\to u}$ all have the same (positive) measure, which we denote by $m$.
Corollary 4.5 (applied with $L=L^{0}$ and $K=\mathbb{D}$, so that $\mathrm{Aut}(K)=\mathrm{Bij}(\mathbb{D})$) ensures that the $\mathsf{H}$-stabilizer of any element $u\in\mathbb{D}$ is measure equivalent to $\mathsf{G}_{u}$, whence finite. Let $k$ be a bound on the cardinality of a finite subgroup of $\mathsf{H}$. For every $v\in\mathsf{G}\cdot u$, the set $\Omega_{u\to v}$ is $\mathsf{H}_{u}$-invariant; we let $\Omega^{\prime}_{u\to v}\subseteq\Omega_{u\to v}$ be a fundamental domain for the action of the finite group $\mathsf{H}_{u}$. Then $\Omega^{\prime}_{u\to v}$ has measure at least $m/k$.
One has $\Omega=\coprod_{u\in\mathbb{D}}\Omega_{u\to v}$. Therefore, if $T\subseteq\mathbb{D}$ is a set of representatives of the $\mathsf{H}$-orbits, then $\coprod_{u\in T}\Omega^{\prime}_{u\to v}$ is a fundamental domain for the $\mathsf{H}$-action on $\Omega$. Since $\mathsf{H}$ has a finite measure fundamental domain, and the measure of the above fundamental domain is at least equal to $m|T|/k$, it follows that $|T|<+\infty$. This concludes our proof. ∎
@@
Let $\mathsf{G}$ be a countable group, acting faithfully by automorphisms on a polyhedral complex $K$ with countably many cells, in such a way that $\mathsf{G}$ and $\mathrm{Aut}(K)$ have the same orbits of vertices. Let $\mathsf{H}$ be a countable group with bounded finite subgroups that is measure equivalent to $\mathsf{G}$, let $(\Omega,\mu)$ be a measure equivalence coupling between $\mathsf{G}$ and $\mathsf{H}$, and assume we are given a homomorphism $\iota:\mathsf{H}\to\mathrm{Aut}(K)$ and a measurable almost $(\mathsf{G}\times\mathsf{H})$-equivariant map $\theta:\Omega\to\mathrm{Aut}(K)$.
@@
Proposition 4.7 and Corollary 4.8 fail if one does not assume that $\mathsf{H}$ has bounded finite subgroups. Indeed, it is possible to find a countable group $\mathsf{G}$ acting properly and cocompactly on a polyhedral complex $K$, satisfying the assumptions of Corollary 4.8, and a non-uniform lattice $\mathsf{H}$ in $\mathrm{Aut}(K)$, acting with infinitely many orbits of vertices. In this case $\mathsf{G}$ and $\mathsf{H}$ are measure equivalent.
@@ 4.4 Measure equivalence couplings and index
Let $\mathsf{G}$ and $\mathsf{H}$ be two countable groups, and let $\Sigma$ be a measure equivalence coupling between $\mathsf{G}$ and $\mathsf{H}$. Let $\mathsf{H}^{\prime}\subset\mathsf{H}$ be a subgroup. Assume that there exists a positive measure $(\mathsf{G}\times\mathsf{H}^{\prime})$-invariant Borel subset $\Sigma^{\prime}\subset\Sigma$ such that for any $h\in\mathsf{H}\setminus\mathsf{H}^{\prime}$ we have $\mu(h\Sigma^{\prime}\cap\Sigma^{\prime})=0$.
Then $\mathsf{H}^{\prime}$ is of finite index in $\mathsf{H}$.
@@
Let $E$ and $E^{\prime}$ be respective Borel fundamental domains for the actions of $\mathsf{G}$ and $\mathsf{H}^{\prime}$ on $\Sigma^{\prime}$, chosen such that $U=E\cap E^{\prime}$ has positive measure (these exist because $\mathsf{G}$ and $\mathsf{H}$ admit Borel fundamental domains on $\Sigma$).
We claim that $E$ is contained in a Borel fundamental domain $X_{\mathsf{G}}$ for the $\mathsf{G}$-action on $\Sigma$, and $E^{\prime}$ is contained in a Borel fundamental domain $X_{\mathsf{H}}$ for the $\mathsf{H}$-action on $\Sigma$. For the first part of the claim, we take any Borel fundamental domain $X^{\prime}_{\mathsf{G}}$ for the $\mathsf{G}$-action on $\Sigma$ and take $X_{\mathsf{G}}=(X^{\prime}_{\mathsf{G}}\setminus\Sigma^{\prime})\cup E$. For the second part of the claim, it suffices to prove that for any $h\in\mathsf{H}$, the intersection $hE^{\prime}\cap E^{\prime}$ has measure zero. If $h\in\mathsf{H}^{\prime}$, then this follows from that $E^{\prime}$ is a fundamental domain for the $\mathsf{H}^{\prime}$-action on $\Sigma^{\prime}$. If $h\notin\mathsf{H}^{\prime}$, then our assumption implies that $\mu(h\Sigma^{\prime}\cap\Sigma^{\prime})=0$, and the claim follows.
@@
As recalled in Section 4.1, there is a natural measure-preserving action $\mathsf{H}\curvearrowright X_{\mathsf{G}}$, obtained from the identification $X_{\mathsf{G}}\approx\mathsf{G}\backslash\Sigma$. We let $\mathcal{G}_{1}$ be the associated measured groupoid over $X_{\mathsf{G}}$; it is equipped with a natural cocycle $\mathcal{G}_{1}\to\mathsf{H}$. Similarly, we have a natural measure-preserving action $\mathsf{G}\curvearrowright X_{\mathsf{H}}$, and we let $\mathcal{G}_{2}$ be the associated measured groupoid over $X_{\mathsf{H}}$. It is equipped with a natural cocycle $\mathcal{G}_{2}\to\mathsf{G}$. Let now $U_{1}=X_{\mathsf{G}}\cap X_{\mathsf{H}}$: this is a positive measure Borel subset (because it contains $U=E\cap E^{\prime}$). We have $(\mathcal{G}_{1})_{|U_{1}}=(\mathcal{G}_{2})_{|U_{1}}$. We denote by $\mathcal{G}$ this measured groupoid over $U_{1}$; up to restricting $U_{1}$ to a conull Borel subset, $\mathcal{G}$ is naturally equipped with two cocycles $\rho_{\mathsf{G}}:\mathcal{G}\to\mathsf{G}$ and $\rho_{\mathsf{H}}:\mathcal{G}\to\mathsf{H}$.
We now claim that $\mathcal{G}_{|U}=\rho_{\mathsf{H}}^{-1}(\mathsf{H}^{\prime})_{|U}$. The lemma follows from this claim by using e.g. [HH22b, Lemma B.3].
To prove the claim, it suffices to show that $\mathcal{G}_{|U}\subseteq\rho_{\mathsf{H}}^{-1}(\mathsf{H}^{\prime})_{|U}$. Let $\gamma\in\mathcal{G}_{|U}$, let $u=s(\gamma)$ be the source of $\gamma$. As $\Sigma^{\prime}$ is $\mathsf{G}$-invariant, we have $gu\in\Sigma^{\prime}$. Since $E^{\prime}$ is a fundamental domain for the action of $\mathsf{H}^{\prime}$ on $\Sigma^{\prime}$, there exists a unique $h\in\mathsf{H}^{\prime}$ such that $hgu\in E^{\prime}$. So $h$ is also the unique element of $\mathsf{H}$ satisfying $hgu\in X_{\mathsf{H}}$. Therefore $\rho_{\mathsf{H}}(\gamma)=h$, showing that $\gamma\in\rho_{\mathsf{H}}^{-1}(\mathsf{H}^{\prime})$, as desired. ∎
@@ 5 Proximal dynamics and strong ICC property for $\mathrm{Aut}(\mathbb{B})$
The goal of this section is to prove the following proposition. We refer to Definition 4.1 for the definition of the strong ICC property.
@@
Let $G$ be a right-angled Artin group with trivial center, and let $\mathbb{B}$ be its right-angled building. Then the inclusion $G\subseteq\mathrm{Aut}(\mathbb{B})$ is strongly ICC.
The assumption that $G$ has trivial center amounts to requiring that in the defining graph of $\Gamma$, no vertex is joined to all other vertices by an edge.
Our plan is to embed $\mathrm{Aut}(\mathbb{B})$ measurably into a bigger group, namely the homeomorphism group $\mathrm{Homeo}(R(\mathbb{B}))$ of the regular part of the Roller compactification of $\mathbb{B}$ (defined in Section 5.2), and show that the inclusion $G\subseteq\mathrm{Homeo}(R(\mathbb{B}))$ is strongly ICC using properties of the dynamics of the action of $G$ on the compact topological space $R(\mathbb{B})$, which implies the proposition. In Section 5.1, we give a general dynamical criterion for proving strongly ICC, and Section 5.2, we justify the measureable embedding of $\mathrm{Aut}(\mathbb{B})$ into $\mathrm{Homeo}(R(\mathbb{B}))$, and show the action of $G$ on $R(\mathbb{B})$ satisfies the dynamical criterion.
@@ 5.1 A strong ICC lemma, after Bader–Furman–Sauer
Recall that an action of a countable group $G$ by homeomorphisms on a compact metrizable space $K$ is strongly proximal in the sense of Furstenberg [Fur63] if the $G$-orbit of every $\nu\in\mathrm{Prob}(K)$ contains a Dirac mass in its weak-* closure. The action is minimal if every $G$-orbit is dense in $K$. Given a compact metrizable space $K$, the group $\mathrm{Homeo}(K)$ is equipped with the topology of uniform convergence, which turns it into a Polish group. The following lemma is a small variation over an argument of Bader–Furman–Sauer [BFS13, Lemma 2.4].
@@
be the space of Borel $\mu$-stationary probability measures on $K$, which is a nonempty compact subset of $\mathrm{Prob}(K)$, equipped with the weak-* topology. Note that $\mathrm{Prob}_{\mu}(K)$ is $G$-invariant, using that $\mu$ is conjugation-invariant.
@@
We are thus left with proving the above claim. Let $\nu\in\mathrm{Prob}_{\mu}(K)$. For every $i\in\{1,\dots,k\}$, let $\nu_{i}\in\mathrm{Prob}(K_{i})$ be the pushforward of $\nu$ under the projection map $K\to K_{i}$. Since the $G_{i}$-action on $K_{i}$ is minimal and strongly proximal, we can find a sequence $(g_{i,n})_{n\in\mathbb{N}}\in G_{i}^{\mathbb{N}}$ such that $(g_{i,n})_{\ast}\nu_{i}$ converges to $\delta_{x_{i}}$, as $n$ goes to $\infty$, in the weak-* topology.
For every $n\in\mathbb{N}$, let $g_{n}=(g_{1,n},\dots,g_{k,n})$. We will now prove that the probability measures $(g_{n})_{\ast}\nu$ converge to $\delta_{x}$, as $n$ goes to $\infty$, in the weak-* topology. This will conclude our proof.
@@
as $n$ goes to $\infty$. Intersecting these sets as $i$ varies in $\{1,\dots,k\}$, we obtain that $\nu(g_{n}^{-1}(U_{1}\times\dots\times U_{k}))\to 1$ as $n$ goes to $\infty$. In particular $(g_{n})_{\ast}\nu(U)\to 1$. Since $U$ was an arbitrary open set containing $x$, by the Portmanteau Theorem, this is enough to prove that $(g_{n})_{\ast}\nu$ converges to $\delta_{x}$, as desired. ∎